OpenStudy (anonymous):

How to find the coordinates of a centroid if my median equations are y=-X+2, y=1, and y=2x-1?

OpenStudy (anonymous):

And that's my centroid?

OpenStudy (ranga):

Centroid is where the medians intersect. Solve for (x,y) in any two of the above three equations. Pick y = 1 as one of the equations because it is so easy.

OpenStudy (anonymous):

What do I do after I solve for X and Y? Or is that it?

OpenStudy (ranga):

y=1, and y=2x-1 put y = 1 in y = 2x - 1 1 = 2x - 1 2x = 2 x = 1 So (1, 1) is the centroid. You will notice it satisfies all three equations. But you will need just two equations to solve for it.

OpenStudy (anonymous):

oh, ok. Thanks a ton man! That's all there is to it?

OpenStudy (ranga):

Yep. You are welcome.

OpenStudy (anonymous):

I'll give you a BJ for a job well done

OpenStudy (anonymous):

Alright. Is that you in your avatar?

OpenStudy (anonymous):

Do you have a math degree?